OpenStudy (anonymous):

The point-slope form of the equation of the line that passes through (–4, –3) and (12, 1) is y – 1 = (x – 12). What is the standard form of the equation for this line?

OpenStudy (anonymous):

y-Y=m(x-X) y-1=m(x-12) m=(1+3)/(12+4)=4/16=1/4 y-1=1/4(x-12) y-1=1/4x-3 y=1/4x-2 y=mx+c standard form